Three times the smaller of two consecutive even integers is the same as -146 minus 4 times the larger integer. Find the integers.

Answer by algebrahouse.com(1659)

x = first consecutive even integer
x + 2 = second consecutive even integer

3x = -146 - 4(x + 2) {three times smaller equals -146 - 4 times larger}
3x = -146 - 4x - 8 {used distributive property}
7x = -154 {added 4x to each side and combined like terms}
x = -22 {divided each side by 7}
x + 2 = -20 {substituted -22, in for x, into x + 2}

-22 and -20 are the two consecutive even integers

For more help from me, visit: www.algebrahouse.com